A certain type of computer costs $1,055 and the annual holding cost of each computer is 24.6 % of the cost. Annual demand is 9,423 units and the ordering cost is $164.4 per order. Calculate the EOQ (economic order quantity) rounded to 1-place of decimal.
Answer #1

DEMAND = 9423
ORDERING COST = 164.4
HOLDING COST = 259.53(24.6% of 1055)
COST PER UNIT = 1055

EOQ = SQRT(2 * DEMAND * ORDERING COST / HOLDING COST) = SQRT(2 * 9423 * 164.4 / 259.53) = 109.3